OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -3, an = 2 ● an-1

OpenStudy (ybarrap):

Is this the problem? See attached

OpenStudy (anonymous):

yes

OpenStudy (ybarrap):

a1 = -3 a2=(2 * -3) -1=-6-1=-7 a3=(2*-7)-1=-14-1=-15 a4=(2*-15)-1=-30-1=-31 a5 = ... a6 = ... Using an = 2 ● an-1. That's it!

OpenStudy (anonymous):

thank you so much !

OpenStudy (ybarrap):

NP